Hair texture in humans is incompletely dominant. Curly hair (C) plus straight hair (c) gives you wavy hair. Also, having 2 eyebrows is dominant over one eyebrow. If two individuals are heterozygous for both traits, what is the probability of each phenotype in their future offspring?
Answer #1

alleles are C-curly hair, c- straight hair, T-two eye brows, t- one eye brow.

CcTt * CcTt

CT Ct cT ct
CT CCTT (curly hair, two eyebrows) CCTt(curly hair, two eyebrows) CcTT( wavy hair, two eyebrows) CcTt( wavy hair, two eyebrows)
Ct CCTt(curly hair, two eyebrows) CCtt ( curly hair, one eyebrow) CcTt( wavy hair, two eyebrows) Cctt ( wavy hair, one eyebrow)
cT CcTT ( wavy hair, two eyebrows) CcTt( wavy hair, two eyebrows) ccTT ( straight hair,2 eyebrows) ccTt( straight hair,2 eyebrows)
ct CcTt( wavy hair, two eyebrows) Cctt ( wavy hair, one eyebrow) ccTt( straight hair,2 eyebrows) cctt ( straight hair, one eyebrow)

possible phenotypes in the progeny are given in the punnet square.

probability of getting curly hair, two eyebrows= number of curly hair, two eyebrows/total number=3/16

probability of getting wavy hair, two eyebrows = number of wavy hair, two eyebrows/total number=6/16

probability of getting curly hair, one eyebrow = number of curly hair, one eyebrow/total number=1/16

probability of getting wavy hair, one eyebrow = number of wavy hair, one eyebrow/total number=2/16

probability of getting straight hair,2 eyebrows= number of straight hair,2 eyebrows/total number=3/16

probbaility of getting straight hair, one eyebrow= number of straight hair, one eyebrow/total number=1/16